A certain artifact was found in 2020 to contain presently 83% of its original carbon-14. In what year did the artifact first exist? The half-life of carbon-14 is approximate 5,740 years.

1
Expert's answer
2021-12-17T13:33:58-0500

The amount of time "t" passed from the moment the artifact first existed can be found from the radioactive decay law:


"N = N_0\cdot 2^{-\dfrac{t}{T}}"

where "N\/N_0 = 0.83", "T = 5740years". Expressing "t", obtain:


"t = -T\log_2\dfrac{N}{N_0} = -5740\cdot \log_20.83 \approx 1543years"

Since now is year 2020, the artifact first existed in year "2020 - 1543 = 477"

Answer. 477.